If a four wheeler travels 7 miles in one hour 29 minutes and 39 seconds. How fast would the four wheeler be traveling?

Answer by Alan3354(69443)

If a four wheeler travels 7 miles in one hour 29 minutes and 39 seconds. How fast would the four wheeler be traveling?
-------------
29' 39" = 29*60 + 39 = 1779 seconds
-----
d = r*t
r = d/t = 7/1779 mi/sec
=~ 0.003935 miles/second